if a car is traveling at a constant rate of 45 miles per hour, how many miles does it travel from 10:40 A.M. to 1:00 P.M. of the same day?
Found 2 solutions by rfer, solver91311:
Answer by rfer(12644)

45*2 1/3=
45*7/3=315/3=105 miles

Answer by solver91311(16868)



20 minutes is one-third of an hour. So from 10:40 AM to 1:00 PM of the same day is 2 and one-third hours. Multiply:



John